# Tidally Heated Sub-Neptunes, Refined Planetary Compositions, and Confirmation of a Third Planet in the TOI-1266 System

### v1

@@ 1 Introduction
But it is unclear whether or not these few planets are representative of the true diversity in the M dwarf sub-Neptune population. Unfortunately, there are currently only a handful of temperate sub-Neptune-sized planets orbiting M dwarfs with well-measured masses and radii that are also favorable targets for atmospheric characterization with JWST. There are currently only seven planets with radii between and K with scheduled JWST observations (TOI-270 d, LP 791-18 c, LTT 3780 c, K2-18 b, TOI-776 c, TOI-1468 c, and TOI-4336 b).
@@
Part of the reason why so few of these planets have scheduled JWST observations is the lack of good mass measurements. Measuring masses of temperate sub-Neptunes with radial velocities (RVs) can be challenging because the host stars are faint and/or active, and low-mass planets with long orbital periods have correspondingly small RV semi-amplitudes. For planets in dynamically interacting multi-planet systems, the masses can be measured with transit timing variations (TTVs, e.g. Greklek-McKeon et al., 2023). The size of the TTV signal is maximized in systems where the planets have orbital period ratios close to mean-motion resonance (MMR; Lithwick et al., 2012), and in systems with larger planet-star mass ratios (e.g., M dwarfs; Agol et al., 2021). The Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) has detected many such systems orbiting M dwarfs and obtained an initial set of transit timing measurements. However, the precision of TTV-based mass estimates depends on the orbital architecture of the system, the sampling of the TTV oscillations, and on our ability to obtain precise transit mid-times. Because of this, planets that are observed at low SNR with TESS require additional high SNR transit observations at later epochs in order to obtain useful mass constraints (Ballard, 2019).
@@
In this study we focus on TOI-1266, a nearby ( pc) M3 star ( K, Stefánsson et al., 2020) that hosts two sub-Neptune-sized transiting planets with equilibrium temperatures of 415 K and 346 K, respectively. Systems with multiple planets of different sizes and bulk densities such as TOI-1266 are uniquely advantageous for testing theories of planet formation and evolution (e.g., Diamond-Lowe et al., 2022; Piaulet et al., 2023) because the planets formed from the same protoplanetary disk and experienced the same irradiation history scaled to their orbital separations. We can use these multi-planet systems to explore how small differences in initial conditions and irradiation environment can lead to different final planetary composition types.
@@
The TOI-1266 system has been previously studied in the literature, with two discovery papers characterizing the properties of the inner two transiting planets (Demory et al., 2020; Stefánsson et al., 2020, hereafter D20 and S20), followed by a paper presenting detailed atmospheric modeling for TOI-1266 c (Harman et al., 2022). This system was also the subject of a separate RV analysis that measured the masses of both transiting planets and reported tentative evidence for a third non-transiting planet candidate (Cloutier et al., 2024, hereafter C24). These three studies find that the inner two TOI-1266 planets ( = 2.59 0.10 , = 2.04 0.11 , = 4.40 , = 3.12 0.76 , C24) display an ‘inverted’ architecture with a larger interior planet. Most multi-planet systems that span the radius valley have the smaller super-Earth on an interior orbit to the larger sub-Neptune (e.g., Weiss et al., 2018). The TOI-1266 system therefore runs counter to the canonical narrative that it should be easier for more distant planets to retain larger primordial hydrogen-rich envelopes, even when accounting for the lower mass of TOI-1266 c (C24). Notably, TOI-1266 b has a higher Transmission Spectroscopy Metric (TSM; this indicates the planet’s relative favorability for atmospheric characterization with transmission spectroscopy, see Kempton et al., 2018) than all currently selected JWST sub-Neptune ( ) targets cooler than 500 K aside from L98-59 d.
@@
In this paper, we significantly extend the TTV baseline with new space- and ground-based observations, and carry out a joint TTV and RV analysis of TESS photometry including new data from the second extended mission, new ground-based photometric follow-up from the Wide-field InfraRed Camera (WIRC) at Palomar Observatory, the Multicolor Simultaneous Camera for studying Atmospheres of Transiting exoplanets (MuSCAT) instrument series at the National Astronomical Observatory of Japan (NAOJ) and the Las Cumbres Observatory Global Telescope network (LCOGT), a Sinistro imager at LCOGT, and archival RV measurements from the High Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N).
@@ 3 Stellar Characterization
We perform an updated analysis to improve the iron abundance constraint, and measure several other refractory element stellar abundances including [Mg/H] and [Si/H] for the first time. The detailed abundance analysis was carried out with an implementation of The Cannon (Behmard et al. in prep.), a data-driven method capable of inferring stellar abundances that does not rely on stellar evolution models. This makes The Cannon an excellent choice for characterizing M dwarfs, which have notoriously complex spectra due to the presence of molecular features. This Cannon implementation was trained on M dwarfs with FGK companions from the Sloan Digital Sky Survey-V/Apache Point Observatory Galactic Evolution Experiment (SDSS-V/APOGEE). SDSS-V/APOGEE is a high resolution ( 22,500), -band (1.511.7 m) spectroscopic survey. Its wavelength coverage is ideal for M dwarfs, which have peak brightness in the near-infrared.
@@ 4.1 TESS
We constructed our TESS photometric model using the exoplanet package (Foreman-Mackey et al., 2021) for the transit light curve component. Our transit model includes the following free parameters: the stellar radius , impact parameter , scaled semimajor axis , and individual transit mid-center times. We used separate planet-to-star radius ratios / for each segment of TESS data (separated by PM, EM, and SEM) to account for possible changes in transit depth (see §4.4). We allowed the mid-time of each transit to vary individually in our fit using the TTVOrbit module of exoplanet, rather than fitting for a constant orbital period and corresponding to a linear ephemeris.
@@
We fixed the planetary eccentricities to zero in the transit fit, as our TTV modeling indicates that both planets must have orbital eccentricities less than 0.1 (see §6.2) and the corresponding effect on the transit light curve shape is therefore negligible. We also fixed the quadratic limb darkening parameters = 0.3442 and = 0.2006. These are the values predicted by the ldtk (Parviainen & Aigrain, 2015) package using the stellar temperature, [Fe/H], and reported in Table 2. We list the priors for our transit model and GP hyperparameters in Table 3. Although we used uniform priors for most parameters, we placed a Gaussian prior on and based on the stellar mass, radius, and planetary orbital period from S20. We used the PyMC3 package (Salvatier et al., 2016) to sample the posterior distribution of our model with No U-Turn Sampling (NUTS), with four parallel chains run with 5000 burn-in steps and 3000 posterior sample draws. We confirmed that the chains evolved until the Gelman-Rubin statistic values are < 1.01 for all parameters. The posterior distributions of each parameter are summarized in Table 3. The detrended transit light curves and their best fit models are shown in Figure 1.
@@ 4.3 MuSCAT imagers and Sinistro
The MuSCAT and MuSCAT3 observations include simultaneous photometry in three or four bands, respectively, and each night has a different transit coverage fraction. We initially allowed / to vary for each band and each night, fitting for individual values shared across all bands for a given night, with all other transit shape parameters shared across bands and nights. We fit the MuSCATs + Sinistro transits jointly with the phased TESS transit profile, used the same priors on the transit parameters as in our WIRC modeling, and fixed the quadratic limb darkening parameters to the predicted values from ldtk for each bandpass. For , we use = 0.5976 and = 0.1739, for we use = 0.5674 and = 0.1542, for we use = 0.3900 and = 0.1957, and for we use = 0.3064 and = 0.1924. We explored the parameter space with the NUTS sampler in PyMC3 for 2000 tune and 2000 draw steps and verified that the Gelman-Rubin statistic values are for all parameters. We then determined the significance of the / constraint in each band on each night, and excluded any light curves with a detection significance below 3 as a metric to include only significant transit detections in our transit depth/timing analysis. This eliminated most partial transit observations and some light curves in or band with higher noise levels. We then repeated the same fitting procedure with a single / parameter shared across bands for a given night in order to account for possible changes in transit depth (see §4.4) and to obtain a final set of transit mid-times. As before, we confirmed that all transit mid-times had normally distributed posterior distributions. Our measured transit times are provided in Table 4 in the Appendix, and the final transit light curves for TOI-1266 b and c are shown in Figures 3 and 4, respectively.
@@ 5 TTV Modeling
The addition of the TESS SEM data nearly doubles our observational baseline and allows us to detect TTVs for planet b at high significance. The main oscillation period of this TTV signal is approximately twice the baseline of the data examined in previous studies and still only a small fraction of the predicted TTV super-period ( = 36 yr). This explains the previous non-detections: With data covering only a small fraction of the TTV super-period, this signal can be effectively removed by changes to the linear ephemeris of planet b in the TTV retrieval. In the following sections, we outline our analysis of these TTVs, which exhibit a long-term trend for TOI-1266 b that is driven by near-resonant commensurability with a third external planet, and short-term chopping signatures for both TOI-1266 b and c. The low-significance TTV detection reported in D20 is likely the same second-order short-term chopping signature we detect here. D20 ascribed this signal to near-resonant 2-planet interactions between TOI-1266 b and c, but we now understand it to be part of a more complex 3-planet system in which the inner and outer planets are near the 3:1 mean-motion commensurability.
@@ 5.1 The 2-Planet Case
In our TTV modeling, we fixed the planetary orbital inclinations () to 90∘ because our transit fits show a low mutual inclination between planets b and c and are very close to edge-on ( = 89.12 0.08, = 89.17 0.04). The difference between and sin() is 0.01% for both transiting planets. The TTV solution is second-order in mutual inclination, with the strength of the planet-planet gravitational interaction and corresponding TTV amplitudes diminishing rapidly as the planetary mutual inclination increases (Nesvorný & Vokrouhlický, 2014; Hadden & Lithwick, 2016). For a purely edge-on orbital inclination the longitude of the ascending node () becomes undefined, so we arbitrarily set it to 90∘ for both planets.
Our 2-planet TTV model has ten free parameters in total. These include: the planet-to-star mass ratios, Keplerian orbital periods, mean anomalies reparamaterized with the time of first transit (), and the planetary eccentricities and longitudes of periastron. We reparameterized the latter two quantities as and in order to mitigate the degeneracy between e and in our fits while retaining an effective uniform prior on (Eastman et al., 2013). The orbital periods, mean anomalies, eccentricities, and longitudes of periastron are osculating orbital elements defined at the TTV model start time . We fit this model to the data using the affine invariant Markov chain Monte Carlo (MCMC) ensemble sampler emcee (Foreman-Mackey et al., 2013), and chose wide uniform priors for all parameters: U(-1, 1) for and , U(0, 30) for the planetary masses, and U(-100, +100) for the planetary orbital period and values using the values reported in C24.
@@
We initially ran the sampler for 10000 steps in order to allow the 2000 walkers randomly distributed across the prior volume to identify the highest likelihood regions of posterior space. We then sorted the walkers by increasing likelihood and retained the walkers with median likelihoods in the top 20% from the last 100 steps. Finally, we ran the remaining walkers for steps. We confirmed that the length of the MCMC chain was at least 10 autocorrelation lengths for all parameters. We plot 100 random draws from the resulting TTV posterior distribution and compare them to the observed TTVs in the top panel of Figure 6.
The 2-planet TTV retrieval identifies two distinct families of solutions – neither of which are a satisfactory fit to the data. One family of solutions prefers very low mass values for both planets, with 95% upper limits of and , and high eccentricity values, and . This set of models (shown in purple in Figure 6) is better at reproducing the long-term TTV variability revealed in the TESS SEM data, but does not accurately model the TESS PM data and struggles to reproduce the short-timescale chopping TTV signature. Unlike the dominant long-term TTV trend with a super-period defined by the proximity of the planetary orbital periods to resonance (Lithwick et al., 2012), the short-timescale chopping signature super-imposed on this trend is defined by the planetary mass and eccentricity vectors (Deck & Agol, 2015). The other family of solutions (shown in green in Figure 6) prefers lower planetary eccentricities and and a very large mass for TOI-1266 c () that is inconsistent with the RV mass constraint reported in C24, and cannot reproduce the long-term TTV trend revealed by the TESS SEM data.
@@ 5.2 The 3-Planet Case
We ran the MCMC retrieval utilizing the same general approach as in §5.1. For this fit, we found that fewer walkers were required to identify the highest likelihood region of the parameter space. We therefore used 600 walkers and ran them for a total of steps. We then selected the most likely 50% of walkers and ran them for additional steps. We verified that the number of steps in the resulting chain was at least 10 times the shortest parameter autocorrelation length. The posterior from this blind search strongly favors a planet near the 3:1 commensurability with the inner planet TOI-1266 b, corresponding to an orbital period of 32.509. The BIC value for the 3-planet TTV model compared to a 2-planet model is -10, indicating very strong preference for the 3-planet case (Raftery, 1995).
Results from our 3-planet TTV retrieval are shown in the bottom two panels of Figure 6. The new 3-planet model is capable of simultaneously describing both the long-term TTV oscillation that samples part of the super-period of TOI-1266 b, and the short-term chopping signal for planets b and c. The fitted orbital period for TOI-1266 d from our TTV modeling is larger than the reported value from C24 by 3.2 (see §6.2 for a discussion on the source of this discrepancy), although our TTV model is consistent with the RV modeling of C24 for the planetary masses and eccentricities. We summarize the prior and posterior distributions for all planets from our 3-planet TTV modeling, the RV modeling described in §6.1, and the TTV+RV joint modeling described in §6.2 and in Table 3.
For completeness, we also consider a scenario in which the third planet has an orbital period intermediate between that of TOI-1266 b and TOI-1266 c. We fit this model to the data using the same brute-force TTV retrieval as above, but find that there are no combinations of parameters that can adequately describe the complete set of observations – similar to the results of the 2-planet model fit in §5.1. Lastly, we consider a case in which the third planet is located interior to TOI-1266 b instead of exterior to TOI-1266 c. We find that the posterior distribution for this fit also prefers a third planet near the 3:1 commensurability with TOI-1266 b, corresponding to an orbital period near 3.6 days. The overall quality of this fit is similar to that of the external 3:1 scenario shown in the bottom two panels of Figure 6.

### v3

@@ 1 Introduction
It is unclear whether or not these few planets are representative of the true diversity in the M dwarf sub-Neptune population. Unfortunately, there are currently only a handful of temperate sub-Neptune-sized planets orbiting M dwarfs with well-measured masses and radii that are also favorable targets for atmospheric characterization with JWST. As a result, there are only seven planets with radii between and K with scheduled JWST observations (TOI-270 d, LP 791-18 c, LTT 3780 c, K2-18 b, TOI-776 c, TOI-1468 c, and TOI-4336 b). The predicted amplitude of the atmospheric absorption signal during transit is inversely proportional to the planet’s surface gravity, and well-constrained mass and radius measurements are therefore an essential requirement for JWST observation planning (e.g., Batalha et al., 2017, 2019; Di Maio et al., 2023). For sub-Neptunes, bulk density measurements can also be used in conjunction with transmission spectroscopy to constrain the composition of the planet’s envelope (e.g., Benneke et al., 2024; Piaulet-Ghorayeb et al., 2024).
@@
Measuring masses of temperate sub-Neptunes with radial velocities (RVs) can be challenging because the host stars are faint and/or active, and low-mass planets with long orbital periods have correspondingly small RV semi-amplitudes. For planets in dynamically interacting multi-planet systems, the masses can be measured with transit timing variations (TTVs, e.g. Greklek-McKeon et al., 2023). The size of the TTV signal is maximized in systems where the planets have orbital period ratios close to mean-motion resonance (MMR; Lithwick et al., 2012), and in systems with larger planet-star mass ratios (e.g., M dwarfs; Agol et al., 2021). The Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) has detected many such systems orbiting M dwarfs and obtained an initial set of transit timing measurements. However, the precision of TTV-based mass estimates depends on the orbital architecture of the system, the sampling of the TTV oscillations, and on our ability to obtain precise transit mid-times. Because of this, planets that are observed at low SNR with TESS require additional high SNR transit observations at later epochs in order to obtain useful mass constraints (Ballard, 2019).
@@
In this study we focus on TOI-1266, a nearby ( pc) M3 star ( K, Stefánsson et al., 2020) that hosts two sub-Neptune-sized transiting planets with equilibrium temperatures of 415 K and 346 K, respectively (Demory et al., 2020; Stefánsson et al., 2020, hereafter D20 and S20). Systems with multiple planets of different sizes and bulk densities such as TOI-1266 are uniquely advantageous for testing theories of planet formation and evolution (e.g., Diamond-Lowe et al., 2022; Piaulet et al., 2023) because the planets formed from the same protoplanetary disk and experienced the same irradiation history scaled to their orbital separations. We can use these multi-planet systems to explore how small differences in initial conditions and irradiation environment can lead to different final planetary composition types.
@@
The TOI-1266 system has been previously studied in the literature, with two discovery papers characterizing the properties of the inner two transiting planets (D20 and S20), followed by a paper presenting detailed atmospheric modeling for TOI-1266 c (Harman et al., 2022). This system was also the subject of a separate RV analysis that measured the masses of both transiting planets and reported tentative evidence for a third non-transiting planet candidate (Cloutier et al., 2024, hereafter C24). These three studies find that the inner two TOI-1266 planets ( = 2.59 0.10 , = 2.04 0.11 , = 4.40 , = 3.12 0.76 , C24) display an ‘inverted’ architecture with a larger interior planet. Most multi-planet systems that span the radius valley have the smaller super-Earth on an interior orbit to the larger sub-Neptune (e.g., Weiss et al., 2018). The TOI-1266 system therefore runs counter to the canonical narrative that it should be easier for more distant planets to retain larger primordial hydrogen-rich envelopes, even when accounting for the lower mass of TOI-1266 c (C24). Notably, TOI-1266 b has a higher Transmission Spectroscopy Metric (TSM; this indicates the planet’s relative favorability for atmospheric characterization with transmission spectroscopy, see Kempton et al., 2018, C24, and Table 3) than all currently selected JWST sub-Neptune ( ) targets cooler than 500 K aside from L98-59 d.
@@
In this paper, we extend the TTV baseline by 730 days with new space- and ground-based observations, and carry out a joint TTV and RV analysis of TESS photometry including new data from the second extended mission, new ground-based photometric follow-up from the Wide-field InfraRed Camera (WIRC) at Palomar Observatory, the Multicolor Simultaneous Camera for studying Atmospheres of Transiting exoplanets (MuSCAT) instrument series at the National Astronomical Observatory of Japan (NAOJ) and the Las Cumbres Observatory Global Telescope network (LCOGT), a Sinistro imager at LCOGT, and archival RV measurements from the High Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N).
@@ 3 Stellar Characterization
We perform an updated analysis to improve the iron abundance constraint, and measure several other refractory element stellar abundances including [Mg/H] and [Si/H] for the first time. The detailed abundance analysis was carried out with an implementation of The Cannon (Behmard et al., 2025), a data-driven method capable of inferring stellar abundances that does not rely on stellar evolution models. This makes The Cannon an excellent choice for characterizing M dwarfs, which have notoriously complex spectra due to the presence of molecular features. This Cannon implementation was trained on M dwarfs with FGK companions from the Sloan Digital Sky Survey-V/Apache Point Observatory Galactic Evolution Experiment (SDSS-V/APOGEE). SDSS-V/APOGEE is a high resolution ( 22,500), -band (1.511.7 m) spectroscopic survey. Its wavelength coverage is ideal for M dwarfs, which have peak brightness in the near-infrared.
@@ 4.1 TESS
We constructed our TESS photometric model using the exoplanet package (Foreman-Mackey et al., 2021) for the transit light curve component. Our transit model includes the following free parameters: the stellar radius , impact parameter , scaled semimajor axis , and individual transit mid-center times. We used separate planet-to-star radius ratios / for each segment of TESS data (separated by PM, EM, and SEM) to account for possible changes in transit depth (see §4.4). In §4.2, §4.3, and §4.4, we refer to b, , and / as “transit shape parameters”. We allowed the mid-time of each transit to vary individually in our fit using the TTVOrbit module of exoplanet, rather than fitting for a constant orbital period and corresponding to a linear ephemeris.
@@
We fixed the planetary eccentricities to zero in the transit fit, as our TTV modeling indicates that both planets must have orbital eccentricities less than 0.1 (see §6.2) and the corresponding effect on the transit light curve shape is therefore negligible. We also fixed the quadratic limb darkening parameters = 0.3442 and = 0.2006. These are the values predicted by the ldtk (Parviainen & Aigrain, 2015) package using the stellar temperature, [Fe/H], and reported in Table 2. ldtk uses the library of high-resolution synthetic spectra from Husser et al. (2013), which are based on the PHOENIX stellar atmosphere code and includes synthetic spectra for stellar temperatures as low as 2300 K. We list the priors for our transit model and GP hyperparameters in Table 3. Although we used uniform priors for most parameters, we placed a Gaussian prior on and based on the stellar mass, radius, and planetary orbital period from S20. We used the PyMC3 package (Salvatier et al., 2016) to sample the posterior distribution of our model with No U-Turn Sampling (NUTS), with four parallel chains run with 5000 burn-in steps and 3000 posterior sample draws. We confirmed that the chains evolved until the Gelman-Rubin statistic values are < 1.01 for all parameters. The posterior distributions of each parameter are summarized in Table 3. The detrended transit light curves and their best fit models are shown in Figure 1.
@@ 4.3 MuSCAT imagers and Sinistro
The MuSCAT and MuSCAT3 observations include simultaneous photometry in three or four bands, respectively, and each night has a different transit coverage fraction. We initially allowed / to vary for each band and each night, fitting for individual values shared across all bands for a given night, with all other transit shape parameters shared across bands and nights. We fit the MuSCATs + Sinistro transits jointly with the phased TESS transit profile, used the same priors on the transit parameters as in our WIRC modeling, and fixed the quadratic limb darkening parameters to the predicted values from ldtk for each bandpass. For , we use = 0.5976 and = 0.1739, for we use = 0.5674 and = 0.1542, for we use = 0.3900 and = 0.1957, and for we use = 0.3064 and = 0.1924. We explored the parameter space with the NUTS sampler in PyMC3 for 2000 tune and 2000 draw steps and verified that the Gelman-Rubin statistic values are for all parameters. We then determined the detection significance of the / constraint in each band on each night, and excluded any light curves where / is consistent with 0 within 3. This ensures that we only include strongly detected transits in our subsequent analysis. This eliminated most partial transit observations and some light curves in or band with higher noise levels. We then repeated the same fitting procedure with a single / parameter shared across bands for a given night in order to account for possible changes in transit depth (see §4.4) and to obtain a final set of transit mid-times. As before, we confirmed that all transit mid-times had normally distributed posterior distributions. Our measured transit times are provided in Table 4 in the Appendix, and the final transit light curves for TOI-1266 b and c are shown in Figures 3 and 4, respectively.
@@ 5 TTV Modeling
The addition of the TESS SEM data nearly doubles our observational baseline and allows us to detect TTVs for planet b at high significance. The main oscillation period of this TTV signal is approximately twice the baseline of the data examined in previous studies and still only a fraction of the predicted TTV super-period for TOI-1266 b and d ( 2000 d, see §5.2 and Equation 5 of Lithwick et al., 2012). This explains the previous non-detections: with data covering less than half of the TTV super-period, this signal can be effectively removed by changes to the linear ephemeris of planet b in the TTV retrieval.
In the following sections we outline our analysis of these TTVs, which exhibit a long-term trend for TOI-1266 b that is driven by near-resonant commensurability with a third external planet, and short-term chopping signatures for both TOI-1266 b and c. The low-significance TTV detection reported in D20 is likely the same second-order short-term chopping signature we detect here. D20 ascribed this signal to near-resonant 2-planet interactions between TOI-1266 b and c, but we now understand it to be part of a more complex 3-planet system in which the inner and outer planets are near the 3:1 mean-motion commensurability.
@@ 5.1 The 2-Planet Case
In our TTV modeling, we fixed the planetary orbital inclinations () to 90∘ because our transit fits show a low mutual inclination between planets b and c and are very close to edge-on ( = 89.12 0.08, = 89.17 0.04). The difference between and sin() is 0.01% for both transiting planets. The TTV solution is second-order in mutual inclination, with the strength of the planet-planet gravitational interaction and corresponding TTV amplitudes diminishing rapidly as the planetary mutual inclination increases (Nesvorný & Vokrouhlický, 2014a; Hadden & Lithwick, 2016). For a purely edge-on orbital inclination the longitude of the ascending node () becomes undefined, so we arbitrarily set it to 90∘ for both planets.
Our 2-planet TTV model has ten free parameters in total. These include: the planet-to-star mass ratios, Keplerian orbital periods, mean anomalies reparamaterized with the time of first transit (), and the planetary eccentricities and longitudes of periastron. We reparameterized the latter two quantities as and in order to mitigate the degeneracy between e and in our fits while retaining an effective uniform prior on (Eastman et al., 2013). The orbital periods, mean anomalies, eccentricities, and longitudes of periastron are osculating orbital elements defined at the TTV model start time . We fit this model to the data using the affine invariant Markov chain Monte Carlo (MCMC) ensemble sampler emcee (Foreman-Mackey et al., 2013), and chose wide uniform priors for all parameters: U(-1, 1) for and , U(0, 15) for the planetary masses sampled in space, U(-0.1, +0.1) and U(-0.2, +0.2), and U(-1.0, -1.0) and U(-2.0, -2.0) from the planetary orbital period and values reported in C24.
@@
We initially ran the sampler for 10000 steps in order to allow the 2000 walkers randomly distributed across the prior volume to identify the highest likelihood regions of posterior space. We then sorted the walkers by increasing likelihood and retained the walkers with median likelihoods in the top 20% from the last 100 steps. Finally, we ran the remaining walkers for steps. We confirmed that the length of the MCMC chain was at least 10 autocorrelation lengths for all parameters. We plot 100 random draws from the resulting TTV posterior distribution and compare them to the observed TTVs in Panels 1 and 2 of Figure 6.
The 2-planet TTV retrieval identifies two distinct families of solutions – neither of which are a satisfactory fit to the data. One family of solutions prefers very low mass values for both planets, with 95% upper limits of and , and high eccentricity values, and . This set of models (shown in light blue in Figure 6) is better at reproducing the long-term TTV variability revealed in the TESS SEM data, but does not accurately model the TESS PM data and struggles to reproduce the short-timescale chopping TTV signature. Unlike the dominant long-term TTV trend with a super-period defined by the proximity of the planetary orbital periods to resonance (Lithwick et al., 2012), the short-timescale chopping signature super-imposed on this trend is defined by the planetary mass and eccentricity vectors (Deck & Agol, 2015). The other family of solutions (shown in light gray in Figure 6) prefers lower planetary eccentricities and and a very large mass for TOI-1266 c () that is inconsistent with the RV mass constraint reported in C24, and cannot reproduce the long-term TTV trend revealed by the TESS SEM data.
@@ 5.2 The 3-Planet Case
We ran the MCMC retrieval utilizing the same general approach as in §5.1. For this fit, we found that fewer walkers were required to reach more than 10 autocorrelation lengths for all parameters. We therefore used 600 walkers and ran them for a total of steps. We then selected the most likely 50% of walkers and ran them for additional steps. We verified that the number of steps in the resulting chain was at least 10 times the shortest parameter autocorrelation length. The posterior from this blind search strongly favors a planet near the 3:1 commensurability with the inner planet TOI-1266 b, corresponding to an orbital period of 32.509. The BIC value for the 3-planet TTV model compared to a 2-planet model is -10, indicating very strong preference for the 3-planet case (Raftery, 1995).
Results from our 3-planet TTV retrieval are shown in Panels 3 and 4 of Figure 6. The new 3-planet model is capable of simultaneously describing both the long-term TTV oscillation that samples part of the super-period of TOI-1266 b, and the short-term chopping signal for planets b and c. The fitted orbital period for TOI-1266 d from our TTV modeling is larger than the reported value from C24 by 3.2 (see §6.2 for a discussion on the source of this discrepancy), although our TTV model is consistent with the RV modeling of C24 for the planetary masses and eccentricities. We summarize the prior and posterior distributions for all planets from our 3-planet TTV modeling, the RV modeling described in §6.1, and the TTV+RV joint modeling described in §6.2 and in Table 3.
For completeness, we also consider a scenario in which the third planet has an orbital period intermediate between that of TOI-1266 b and TOI-1266 c. We fit this model to the data using the same brute-force TTV retrieval as above, but find that there are no combinations of parameters that can adequately describe the complete set of observations – similar to the results of the 2-planet model fit in §5.1. Lastly, we consider a case in which the third planet is located interior to TOI-1266 b instead of exterior to TOI-1266 c. We find that the posterior distribution for this fit also prefers a third planet near the 3:1 commensurability with TOI-1266 b, corresponding to an orbital period near 3.6 days. The overall quality of this fit is similar to that of the external 3:1 scenario shown in Panels 3 and 4 of Figure 6.